OpenStudy (anonymous):

If the circumference of a circle is 32 in, find the radius. Find the radius of the circle. A. 16 in B. 10.2 in C. 2.54 in D. 5.1 in

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

32 = 2*pi*r r = 32 / 2*pi = 5.1